# Strategic Data Center Load Shifting: Implications for Market Efficiency and Transmission Value

### v2

@@ Abstract
Data center electricity use may reach 12% of U.S. demand by 2030, alongside growing ability to shift workloads geographically in response to prices or carbon signals. We examine the system-level implications of such strategic flexibility using a bilevel two-zone model that couples economic dispatch with consumer cost minimization. Two market failures emerge. First, discontinuous price changes at generator capacity limits can induce flexible consumers to shift load in socially inefficient directions—for example, toward a higher-cost region to trigger a price drop elsewhere. Second, by positioning near capacity boundaries, consumers can eliminate the marginal value of transmission expansion: although shadow prices suggest additional capacity is valuable, consumers reoptimize to offset resulting flow changes, leaving dispatch and costs unchanged. We derive conditions under which these effects arise and show that conventional price signals and planning methods can misrepresent system value in the presence of large flexible loads.
@@ I Introduction
Data center growth is rapidly reshaping electricity demand. In the U.S., data center consumption may more than triple from 2023 to 2030, reaching $\sim$12% of total demand [mckinsey2024]. Because large operators can migrate workloads, these facilities act not only as major loads but also as sources of spatial and temporal flexibility—capabilities already used to reduce costs or emissions of hyperscale computing [google_env_report, Meta_sust_report, ZhengEtal2020_joule, radovanovic2022carbon].
A common assumption is that this flexibility should improve system efficiency: by shifting load toward locations with lower locational marginal prices (LMPs), flexible consumers will provide a similar benefit to the system as transmission—reducing congestion and better utilizing cheap generation [zhang2020flexibility]. This view, however, assumes consumers are price-takers whose decisions do not influence dispatch. As data center loads grow large enough to affect which generators are marginal, decentralized shifting can alter prices and congestion in ways that may not align with system objectives.
@@
    When can decentralized, price-responsive load shifting increase system costs?
@@
Prior work exploring benefits of flexibility often assumes centralized coordination [zhang2020flexibility, knittel2025] or treats consumers as price/emissions-takers [radovanovic2022carbon, SouzaEtal2023], and consequently does not capture such strategic interactions. While recent work identifies counterintuitive outcomes from decentralized shifting [GorkaEtal2024], the mechanisms by which flexible consumers exert market power, and their implications for transmission value, remain underexplored.
We address this gap using a bilevel model in which a flexible consumer chooses spatial load allocation while the system operator performs economic dispatch. In a two-zone system, we derive closed-form conditions under which decentralized incentives align with or diverge from system objectives. Our contributions are:
@@ II Two-Zone Bilevel Model
We model the interaction between a flexible electricity consumer and a system operator as a bilevel optimization problem. The consumer (upper level) chooses how to allocate load across two zones to minimize procurement costs, while the system operator (lower level) performs economic dispatch given the consumer’s allocation decision.
Generation and network structure. Consider a power system with two zones, $A$ and $B$, each containing a merit-order supply stack of generation resources. Let $\mathcal{G}_{A}$ and $\mathcal{G}_{B}$ denote the ordered sets (with respect to marginal cost) of generators at zones $A$ and $B$ respectively. Each generator $g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}$ is characterized by: (i) the capacity constraint: $0\leq p_{g}\leq P_{g}$ (MW), and (ii) constant marginal generation costs: $C_{g}\geq 0$ (\$/MWh). Let $F\geq 0$ denote the total transmission capacity between the two zones, and let $f\in[-F,F]$ denote the power flow from zone $A$ to zone $B$, where positive values indicate AA-to-$B$ flow.
Load allocation. Each zone has two types of loads: (i) Inflexible base load: $b_{A},\,b_{B}\geq 0$ (MWh), which cannot be shifted; (ii) Flexible consumer load: Initially allocated as $x_{A},\,x_{B}\geq 0$ (MWh). Importantly, the flexible consumer is able to shift load between zones. Let $\delta\in[-\alpha,\alpha]$ denote the net load shift from zone $A$ to zone $B$, where $\delta>0$ (resp. $\delta<0$) means net shift from $A$ to $B$ (resp. from $B$ to $A$). Here, $\alpha>0$ denotes the maximum shifting capability. After shifting, the realized flexible loads are $x_{A}-\delta$ at zone $A$ and $x_{B}+\delta$ at zone $B$.
@@ II-A Lower Level: Economic Dispatch
$G(\delta,F):=\min_{\boldsymbol{p},\,f}\quad$ $\sum_{g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}}C_{g}\,p_{g}$ (1a)
@@
The dual variables have standard economic interpretations: $\lambda_{A},\lambda_{B}$ denotes the Locational Marginal Prices (LMPs) at zones $A$ and $B$; $\mu^{+},\mu^{-}$ the shadow price of upper/lower transmission limits. Importantly, LMPs equal the marginal cost of the marginal generator in each zone.
@@ II-B Upper Level: Flexible Consumer Problem
The flexible consumer chooses a load shift $\delta$ to minimize total procurement cost:
@@
$\min_{-\alpha\leq\delta\leq\alpha}\quad P(\delta,F)$ (3)
@@ II-C Objective Alignment
We say the system and consumer objectives are aligned with respect to shift $\delta$ if:
$\Delta P(\delta)<0\quad\Rightarrow\quad\Delta G(\delta)\leq 0,$ (6)
and misaligned otherwise (i.e., if the flexible consumer prefers a shift that increases system cost).
While deliberately stylized, our two-zone framework captures the key mechanisms of price formation and congestion in power systems while remaining analytically tractable. It also reflects two practically relevant settings for hyperscale load shifting. First, setting $F=0$ models two independent electricity markets (e.g., different continents or asynchronous grids) with no physical power transfer, isolating how differences in local supply stacks—particularly the slope of merit order curves—can create misalignment between system and consumer objectives. Second, setting $F>0$ represents two regions within an interconnected market, highlighting how spatially flexible demand interacts with congestion and marginal pricing to distort the perceived value of transmission capacity and alter investment signals.
@@ III-A Basis Changes and Discontinuities
When load shifts are large enough to change marginal generators, the flexible consumer cost function develops discontinuities that yield potential for misalignment. In Proposition 1, we describe how system and flexible consumer costs change when flexible consumers exert market power over prices. Specifically, Proposition 1 relates changes in the system and flexible consumer objectives at basis changes of (1), denoted by $\tau_{1},\,\tau_{2}$.
@@
Consider a positive load shift $\delta>0$ from $A$ to $B$. Letting $p_{g_{A}},\,p_{g_{B}}$ respectively denote the pre-shift dispatch levels of marginal generators $g_{A}$ and $g_{B}$, two cases arise depending on which zone’s marginal generator changes first.
@@
In other words, when load shifts are small enough that the optimal basis of the dispatch problem does not change, system and consumer objectives are perfectly aligned: both benefit linearly from shifting toward the lower-cost region. In this regime, marginal generation costs act as fully welfare-aligned price signals.
Once a basis change occurs, however, Proposition 1 reveals a fundamental asymmetry. The system pays the cost of a newly marginal generator only on its incremental production, whereas the consumer pays the new marginal price on all load at that location. This discrepancy creates discontinuous jumps in the consumer objective at $\delta=\tau_{1}$ or $\delta=\tau_{2}$. As a consequence, flexible consumers may find it profitable to depress prices by strategically shifting load to alter which generators are marginal in zones $A$ and $B$, even when such shifts increase total system cost.
@@ III-B Strategic Positioning and Misalignment
Theorem 1 formalizes the conditions under which the discontinuities identified in Proposition 1 lead to negative externalities, causing decentralized load shifting to diverge from system-optimal behavior.
@@
Consider a positive load shift $\delta=\tau>0$ from $A$ to $B$ that is exactly the minimal shift required to change the optimal basis of (1) (i.e., to induce a change in the marginal generator at either $A$ or $B$). Let $p_{g_{A}}$ and $p_{g_{B}}$ denote the pre-shift dispatch levels of the marginal generators at $A$ and $B$.
The system and consumer objectives are aligned for all such shifts except when all of the following conditions hold:
@@
When these three conditions hold, the flexible consumer reduces their procurement cost at $\delta=\tau$ while system generation cost increases, yielding misalignment.
Theorem 1 can be interpreted as follows. A flexible consumer may be incentivized to shift load against the direction suggested by the LMP difference when one or more of the following conditions hold:
  1. 1.
    Large cost drop at $\boldsymbol{A}$: the difference $(C_{g_{A}}-C_{g_{A-1}})$ is large.
  2. 2.
    Small inter-zone price difference: the gap $(C_{g_{B}}-C_{g_{A}})$ is small.
  3. 3.
    Large allocation at $\boldsymbol{A}$: the quantity $(x_{A}-\tau_{2})$ is large, so the resulting price drop applies to many megawatt-hours.
  4. 4.
    Small shift required: the threshold $\tau_{2}$ is small, making it easy to trigger the marginal generator change.
Under these conditions, local marginal prices fail to encode system-optimal incentives, and decentralized load shifting can increase total system cost.
@@ III-C Illustrative Numerical Example
    $g_{B}^{1}$: 1.2 GW at \$40/MWh
In this configuration, a flexible consumer minimizes procurement cost by shifting half of their load from $A$ to $B$ so that generator $g_{A}^{2}$ is no longer marginal. This reduces the LMP at $A$ from \$25/MWh to \$0/MWh. However, the resulting load shift forces a larger share of energy to be supplied by $g_{B}^{1}$ at \$40/MWh, increasing total system generation costs. This occurs because the cost gap between $g_{A}^{1}$ and $g_{A}^{2}$ is larger than the gap between $g_{A}^{2}$ and $g_{B}^{1}$, creating an incentive for strategic shifting even though it raises system costs.
@@ IV Transmission-Coupled Markets
We now introduce transmission coupling ($F>0$), which creates additional strategic interactions beyond those in Section III. When load is inflexible, the short-run marginal value of transmission—defined by $\frac{d}{dF}G$, or the reduction in cost per unit of added capacity—is conventionally understood to be the shadow price of the binding transmission constraint [hogan1992contract]. With flexible load, however, locational prices influence the spatial allocation of demand, meaning that transmission and load shifting interact. This raises the question of how price-responsive load shifting affects the marginal value of increasing transmission capacity.
@@ IV-A Marginal Value of Transmission Under Flexible Load
$F_{1}:=F_{0}+\min\left\{\Gamma,\,\alpha-\delta_{0}\right\}$
@@
    The marginal value of transmission with respect to the system objective is zero under load shifting:
@@
    The marginal value of transmission with respect to the flexible consumer objective is negative under load shifting:
@@
As a consequence, the dispatch and total system cost remain constant even as transmission capacity expands: the system-level marginal value of transmission is zero. Meanwhile, flexible consumer costs increase, since maintaining $g_{A}$ as the marginal generator in zone $A$ requires allocating more load to the higher-cost zone $B$. Thus, although transmission expansion appears beneficial when evaluated through its shadow price, this apparent benefit is fully offset by endogenous load shifting.
@@ IV-B Illustrative Numerical Example
Across all $F$, the system-optimal load shift is $\delta=0$, which allows full use of cheaper generation at $A$ and avoids dispatching the high-cost unit $g_{B}^{3}$. However, when $F<\Gamma$ (107 MW), flexible consumers can lower their own costs by shifting load to $B$ to avoid dispatching $g_{A}^{3}$, thereby raising the price at $A$ from \$25/MWh to \$60/MWh. This blocks the system operator from serving $B$ with cheaper power from $g_{A}^{3}$, increasing total system cost (see $F=0$ and $F=100$ in Table I).
@@
$\Delta$ Flexible Cost (\$) 5,000-5{,}000 500-500 0
@@
The implications of Theorem 2 are further illustrated in Fig. 3, which compares system and flexible consumer costs as functions of $F$ both with and without load shifting. For $F<\Gamma$, endogenous load shifting exactly offsets the benefits of transmission expansion, yielding no reduction in system cost while raising costs for the flexible consumer. For $F\geq\Gamma$, this offset vanishes and transmission capacity again provides system value.
Figure 3: Comparison of system (left) and flexible consumer costs (right) without load shifting (blue) and with decentralized load shifting (orange) for a range of transmission capacities.
@@ V Conclusion
In this paper, we develop an analytical framework for understanding how price-responsive flexible consumers interact with power system operations through marginal price signals. By deriving closed-form expressions for both private and social objectives, we characterize when decentralized load-shifting behavior aligns with, or diverges from, system-optimal dispatch. We further analyze the marginal value of transmission capacity in the presence of flexible load and show that conventional investment signals—such as transmission shadow prices—may fail to reflect the true cost-reducing value of capacity expansion when flexible consumers adjust their demand endogenously. This highlights the need to account for strategic load response in capacity expansion modeling as well as designing pricing mechanisms and coordination policies for future power systems with large flexible loads.

### v3

@@ Abstract
Data center electricity use may reach 12% of U.S. demand by 2030, alongside growing ability to shift workloads geographically in response to prices or carbon signals. We examine the system-level implications of such strategic flexibility using a bilevel two-zone model that couples economic dispatch with consumer cost minimization. Two market failures emerge. First, discontinuous price changes at generator capacity limits can induce flexible consumers to shift load in socially inefficient directions—for example, toward a higher-cost region to trigger a price drop elsewhere. Second, by positioning near capacity boundaries, consumers can counteract the marginal benefit of transmission expansion: although shadow prices suggest additional capacity is valuable, strategic consumers reoptimize to offset resulting flow changes, leaving dispatch and costs unchanged. We derive conditions under which these effects arise and show that conventional price signals can misrepresent system value in the presence of large spatially flexible loads.
@@ I Introduction
Data center growth is rapidly reshaping electricity demand. In the U.S., data center consumption may more than triple from 2023 to 2030, reaching $\sim$12% of total demand [4]. Because large operators can migrate workloads, these facilities act not only as major loads but also as sources of spatial and temporal flexibility—capabilities already used to reduce costs or emissions of hyperscale computing [2, 8, 13, 9].
A common assumption is that such flexibility should improve system efficiency: by shifting load toward locations with lower locational marginal prices (LMPs), flexible consumers provide benefits similar to transmission—reducing congestion and better utilizing low-cost generation [12]. This view, however, relies on consumers behaving as price-takers. In contrast, empirical studies of electricity markets document that some large participants strategically exert market power, often at a net cost to social welfare [1, 6]. As data center loads grow large enough to influence prices, incentives emerge for strategic load shifting that undermines system efficiency.
@@
    When can strategic load shifting reduce system efficiency?
@@
Prior work exploring the benefits of flexibility often assumes centralized coordination [12, 7] or treats consumers as price- or emissions-takers [9, 11], and therefore does not capture strategic interactions by large flexible loads. While recent work has begun to identify counterintuitive outcomes from decentralized load shifting with respect to emissions [3], the mechanisms by which flexible consumers influence prices and congestion—and the resulting implications for transmission value—remain underexplored.
We address this gap using a bilevel model in which a flexible consumer chooses spatial load allocation while the system operator clears the resulting market. In a two-zone system, we derive closed-form conditions under which a flexible consumer’s incentives align with or diverge from system objectives. Our contributions are:
@@ II Two-Zone Bilevel Model
We model the interaction between a spatially flexible electricity consumer and a system operator as a bilevel optimization problem. The consumer (upper level) chooses how to allocate load across two zones to minimize procurement costs, while the system operator (lower level) clears the market given a baseline system load and the consumer’s allocation decision.
Generation and network structure. Consider a power system with two zones, $A$ and $B$, each containing a merit-order supply stack of generation resources. Each generator $g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}$ has capacity $0\leq p_{g}\leq P_{g}$ (MW) and constant marginal cost $C_{g}\geq 0$ (\$/MWh). Let $F\geq 0$ denote the transmission capacity between the zones, and let $f\in[-F,F]$ denote power flow from $A$ to $B$.
Generators in each zone are ordered by strictly increasing marginal cost. For a given dispatch, let $g_{i}$ denote the marginal generator at zone $i\in\{A,B\}$, with $g_{i}\pm 1$ denoting the adjacent generators in the merit order when they exist. This strict ordering is without loss of generality, as generators with identical marginal costs can be aggregated into a single unit.
Load allocation. Each zone has two types of loads: (i) Inflexible base load: $b_{A},\,b_{B}\geq 0$ (MWh), which cannot be shifted; (ii) Flexible consumer load: Initially allocated as $x_{A},\,x_{B}\geq 0$ (MWh). Importantly, the flexible consumer can shift load between zones. Let $\delta\in[\underline{\delta},\overline{\delta}]$ denote the net load shift from zone $A$ to zone $B$, where $\delta>0$ (resp. $\delta<0$) means net shift from $A$ to $B$ (resp. from $B$ to $A$). The bounds $\underline{\delta}$ and $\overline{\delta}$ capture operational constraints imposed by the flexible consumer (e.g., service constraints). After shifting, the realized flexible loads are $x_{A}-\delta$ at zone $A$ and $x_{B}+\delta$ at zone $B$.
@@ II-A Lower Level: Economic Dispatch
$G(\delta,F):=\min_{p,\,f}\quad$ $\sum_{g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}}C_{g}\,p_{g}$ (1a)
@@
The dual variables have standard economic interpretations: $\lambda_{A},\lambda_{B}$ denote the LMPs at zones $A$ and $B$; $\mu^{+},\mu^{-}$ the shadow price of upper/lower transmission limits. Importantly, LMPs equal the marginal cost of the marginal generator serving each zone.
We assume that all demand is inelastic and must be fully served (i.e., no load shedding) with willingness-to-pay exceeding the highest marginal generation cost. Under this standard assumption, minimizing the system objective $G(\delta,F)$ is equivalent to maximizing social welfare [10].
@@ II-B Upper Level: Flexible Consumer Problem
The flexible consumer chooses a load shift $\delta$ to minimize their own procurement cost; we assume the consumer has perfect information about parameters of the market clearing problem (1).
@@
$\min_{\underline{\delta}\leq\delta\leq\overline{\delta}}\quad P(\delta,F)$ (3)
@@ II-C Objective Alignment
We say the system and consumer objectives are misaligned with respect to shift $\delta$ if:
$\Delta P(\delta)<0\quad\quad\text{and}\quad\quad\Delta G(\delta)>0,$ (6)
i.e., if the flexible consumer prefers a shift that increases system cost, and aligned otherwise.
While deliberately stylized, our two-zone framework captures key mechanisms of price formation and congestion while remaining analytically tractable. Setting $F=0$ models disconnected markets, isolating how differences in local supply stacks can create misalignment between system and consumer objectives, while $F>0$ represents interconnected regions, highlighting how spatially flexible demand interacts with congestion and marginal pricing to distort transmission value and investment signals.
@@ III-A Basis Changes and Discontinuities
When load shifts are large enough to change marginal generators, the flexible consumer cost function develops discontinuities that can lead to misalignment. Proposition 1 characterizes how system and flexible consumer objectives change at basis changes of (1), denoted by $\tau_{1}$ and $\tau_{2}$.
@@
Without loss of generality, consider a positive load shift $\delta>0$ from $A$ to $B$. Let $p_{g_{A}}$ and $p_{g_{B}}$ denote the pre-shift dispatch levels of the marginal generators $g_{A}$ and $g_{B}$, respectively. We assume $p_{g_{A}}\neq P_{g_{B}}-p_{g_{B}}$ so that exactly one marginal generator changes first as $\delta$ increases. Two cases then arise, depending on which zone’s marginal generator changes first.
@@
In other words, when load shifts are small enough that the optimal basis of the dispatch problem does not change, system and consumer objectives are perfectly aligned: both benefit linearly from shifting toward the lower-cost region.
Once a basis change occurs, however, the consumer pays the new marginal price on all load at affected locations while the system only pays incrementally. This yields discontinuities in the consumer objective at $\delta=\tau_{1}$ or $\delta=\tau_{2}$, creating incentives for flexible consumers to depress prices by strategically shifting load to alter which generators are marginal in zones $A$ and $B$, even when such shifts increase total system cost.
@@ III-B Strategic Positioning and Misalignment
Theorem 1 formalizes the conditions under which these discontinuities induce load shifting outcomes that raise system costs.
@@
Without loss of generality, consider a positive load shift $\delta=\tau>0$ from $A$ to $B$ that is exactly the minimal shift required to change the optimal basis of (1) (i.e., to induce a change in the marginal generator at either $A$ or $B$). Let $p_{g_{A}}$ and $p_{g_{B}}$ denote the pre-shift dispatch levels of the marginal generators at $A$ and $B$.
The system and consumer objectives are misaligned with respect to $\delta$ if and only if all of the following conditions hold:
@@
In other words, a flexible consumer is incentivized to shift load against LMP differences, and consequently raise generation costs, when the immediate cost drop from de-committing a marginal generator dominates the incremental cost of shifting load toward a higher-priced region.
@@ III-C Illustrative Numerical Example
    $g_{B}^{1}$: 1,200 MW at \$40/MWh
In this configuration, a flexible consumer minimizes procurement cost by shifting half of their load from $A$ to $B$ so that generator $g_{A}^{2}$ is no longer marginal. This reduces the LMP at $A$ from \$25/MWh to \$0/MWh. However, the resulting load shift forces a larger share of energy to be supplied by $g_{B}^{1}$ at \$40/MWh, increasing total system generation costs. This occurs because the cost gap between $g_{A}^{1}$ and $g_{A}^{2}$ is large relative to the gap between $g_{A}^{2}$ and $g_{B}^{1}$, creating an incentive for strategic shifting even though it raises system costs.
@@ IV Transmission-Coupled Markets
We now introduce transmission coupling ($F>0$), which creates additional strategic interactions beyond those in Section III. When load is inflexible, the short-run marginal value of transmission—defined by $\frac{d}{dF}G$, or the reduction in cost per unit of added capacity—is conventionally understood to be the shadow price of the binding transmission constraint [5]. With flexible load, however, locational prices influence the spatial allocation of demand, meaning that transmission and load shifting are endogenously coupled. This raises the question of how strategic load shifting affects the marginal value of increasing transmission capacity.
@@ IV-A Marginal Value of Transmission Under Flexible Load
$F_{1}:=F_{0}+\min\left\{\Gamma,\,\overline{\delta}-\delta_{0}\right\}$
@@
    The marginal value of transmission w.r.t. the system objective is zero under load shifting:
@@
    The marginal value of transmission w.r.t. the flexible consumer objective is negative under load shifting:
@@
As a consequence, the dispatch and total system cost remain constant even as transmission capacity expands: the system-level marginal value of transmission is zero. Meanwhile, flexible consumer costs increase, since maintaining $g_{A}$ as the marginal generator in zone $A$ requires allocating more load to the higher-cost zone $B$. Thus, although transmission expansion appears beneficial when evaluated through its shadow price, this apparent benefit is fully offset by strategic load shifting.
@@ IV-B Illustrative Numerical Example
For all $F$, the system-optimal load shift is $\delta=0$, which allows full use of cheaper generation at $A$ and avoids dispatching the high-cost unit $g_{B}^{3}$. However, when $F<\Gamma$ (107 MW), flexible consumers reduce their own costs by shifting load to $B$ to avoid dispatching $g_{A}^{3}$, thereby raising the price at $A$ from \$25/MWh to \$60/MWh. This blocks the system operator from serving $B$ with cheaper power from $g_{A}^{3}$, increasing total system cost (see $F=0$ and $F=100$ in Table I).
@@
$\Delta$ Flexible Cost (\$) -5,000 -500 0
@@
The implications of Theorem 2 are further illustrated in Fig. 3, which compares system and flexible consumer costs with respect to $F$ both with and without load shifting. For $F<\Gamma$, strategic load shifting exactly offsets the benefits of transmission expansion, yielding no change in system cost while raising costs for the flexible consumer. For $F\geq\Gamma$, this offset vanishes and transmission capacity again provides system value.
Figure 3: Comparison of system (left) and flexible consumer costs (right) without load shifting (blue) and with strategic load shifting (orange) for a range of transmission capacities.
@@ V Conclusion
We develop an analytical framework showing how price-making behavior by flexible consumers can reduce market efficiency and distort standard investment signals. By deriving closed-form expressions for private and social objectives, we characterize when strategic load shifting aligns with, or diverges from, system-optimal dispatch, and show that conventional measures of transmission value—such as shadow prices—may fail when demand adjusts endogenously. These results underscore the need to account for strategic load response in capacity expansion models and market design.
While our analysis focuses on a two-zone system with perfectly informed consumers for analytical tractability, the underlying mechanism—price discontinuities induced by basis changes in dispatch—extends naturally to general networked systems. In larger networks, flexible loads may similarly position near generator or transmission capacity limits to induce discrete price changes. In practice, such strategic positioning depends on consumers’ ability to anticipate price responses, and imperfect information or forecasting error may therefore limit how frequently such behavior arises. Characterizing the resulting misalignment in meshed networks and under imperfect information is an important direction for future work.